OpenStudy (anonymous):

If x = 4 in this equation, what does y equal? 5x + 8y = 4 A. y = 2 B. y = 4 C. y = -2 D. y = -4

OpenStudy (anonymous):

first, put 4 as x in 5x

OpenStudy (anonymous):

got it?

OpenStudy (ivettef365):

5(4) + 8y = 4 20 + 8y = 4 -20 -20 ------------- 8y = -16 --- --- 8 8 therefore, y = -2